I am having trouble solving these problems. Could you kindly help me.
1. 1/3s+7/9=4/3s
2.In 2010 world grain supply will be 1.8 trilliam metric tons and the supply will be only only 3/4 of world grain demand. What will world grain demand be in 2010?
3. Water boils at 212 F
a) Use the accompanying graph to determine the Celsius temperature at which water boils.
b) Find the Farenheit temperature of hot tap water at 70 C by solving the equation
70=5/9(F-32)
4. 9-2m/7=19
Thank you kindly for helping me.

Answer by stanbon(48545)

1. (1/3)s + 7/9 = (4/3)s
[4/3 - 1/3]s = 7/9
s = 7/9
-------------------------

2.In 2010 world grain supply will be 1.8 trillion metric tons and the supply will be only 3/4 of world grain demand. What will world grain demand be in 2010?
1.8 = (3/4)(world demand)
world demand = (4/3)1.8) = 2.4 trillion metric tons
----------------------
3. Water boils at 212 F
a) Use the accompanying graph to determine the Celsius temperature at which water boils.
Comment: I don't see your graph but I know the boiling temperature of water
is 100 degrees Celsius
-----------------------
b) Find the Farenheit temperature of hot tap water at 70 C by solving the equation
70=(5/9)(F-32)
F-32 = (9/5)70
F-32 = 9*14
F = 126+32 = 160 degrees
---------------------------

4. 9-2m/7=19
-2m/7 = 28
-2m = 196
m = - 98
===============
Cheers,
Stan H.